OpenStudy (anonymous):

Find the coefficient a of the given term in the expansion of of the binomial. Binomial: (x^2+3)^12 Term: ax^10

OpenStudy (ybarrap):

We will use the binomial theorem: http://en.wikipedia.org/wiki/Binomial_theorem $$ \large (x^2+3)=\sum_{i=0}^{12} \binom{12}{2}(x^2)^i\times3^{12-i} $$ From which it it is clear that the \(x^{10^{th}}\) term is $$ \large \binom{12}{5}(x^2)^5\times3^7 $$ That's it! Does this make sense?

OpenStudy (anonymous):

I'm pretty confused on this, I just don't understand it @ybarrap

OpenStudy (ybarrap):

Have you worked with the binomial theorem before?

OpenStudy (anonymous):

No this is my first time @ybarrap

OpenStudy (ybarrap):

Do you know what this means? $$ \large (x^2+3)=\sum_{i=0}^{12} \binom{12}{2}(x^2)^i\times3^{12-i} $$

OpenStudy (anonymous):

Yes @ybarrap

OpenStudy (ybarrap):

which by the way should actually be $$ \large (x^2+3)=\sum_{i=0}^{12} \binom{12}{\color{red}{i}}(x^2)^i\times3^{12-i} $$

OpenStudy (ybarrap):

Ok. So if you look at that carefully, look at the \(x^2\) term. What \(i\) would make \((x^2)^i\) equal to \(x^{10}\)?

OpenStudy (anonymous):

Yes

OpenStudy (ybarrap):

What is the value of \(\large i\) that makes $$ \Large (x^2)^i=x^{10} $$

OpenStudy (ybarrap):

$$ \Large (x^2)^i=x^{2i} $$ Does this help?

OpenStudy (anonymous):

Yes

OpenStudy (ybarrap):

Well then what is \(i\)?

OpenStudy (ybarrap):

$$ \huge{ x^{2i}=x^{10} } $$ What is \(\Huge i\)?

OpenStudy (ybarrap):

$$ \Huge 2i=10 $$ What is \(\huge i\)?

OpenStudy (anonymous):

Would i be 5?

OpenStudy (ybarrap):

Is \(2\times 5=10\)?

OpenStudy (anonymous):

Oh ok I didn;t know you were suppose to multiply

OpenStudy (ybarrap):

The point is that you need to know which of the terms in the sum are the ones associated with \(x^{10}\) and i=5 is the key. Knowing that i=5, then you can just take out of the sum (in the binomial series) the constants associated with i=5: $$ \large{ (x^2+3)=\sum_{i=0}^{12} \binom{12}{\color{red}{i}}(x^2)^i\times3^{12-i}\\ =\binom{12}{0}(x^2)^0\times3^{12-0}+\binom{12}{1}(x^2)^1\times3^{12-1}+\cdots+\\\color{red}{\binom{12}{5}(x^2)^{5}\times3^{12-5}}+\cdots+\binom{12}{12}(x^2)^{12}\times3^{12-12} } $$. The term in red is the one associated with \(x^{10}\) So now all you have to do is calculate: $$ \large \binom{12}{5}\times3^7 $$ Which is the \(a\) in \(ax^{10}\). I hope this makes sense.

OpenStudy (anonymous):

It will thank you

OpenStudy (ybarrap):

What do you get when you calculate? I'll verify for you.

OpenStudy (ybarrap):

Use this if you need help - http://www.wolframalpha.com/input/?i=%28x^2%2B3%29^12

OpenStudy (anonymous):

I got 1732104 I think I calculated it wrong

OpenStudy (ybarrap):

From which you can validate that our solution is correct - http://www.wolframalpha.com/input/?i=%2812+choose+5%29*3^7